OpenStudy (anonymous):

What is the d value of the following arithmetic sequence? 5, 2, -1, -4, -7,… A. 3 B. -3 C. 4 D. -4

hartnn (hartnn):

d = common difference between the consecutive terms d= 2nd term - 1st term = 3rd term - 2nd term =....and so on here, d =2-5 = -1 - 2 = -4 - (-1) = .... ?

OpenStudy (anonymous):

THE ANSWER IS A @hartnn

hartnn (hartnn):

nopes

OpenStudy (anonymous):

Why not ? @hartnn

hartnn (hartnn):

because d = 2-5 = -3

OpenStudy (anonymous):

5-3=2, 2-3=-1, @hartnn

hartnn (hartnn):

right, so the difference will be -3 right ?? if the sequence was 1,4,7,10....the d would have been 3

OpenStudy (anonymous):

okay i get it now @hartnn

hartnn (hartnn):

good :)